# Invariants for $\mathbb{G}_{(r)}$-modules

### v1

@@ Abstract.
Following earlier work of C. Bendel, J. Petvsova, P. Sobaje, A. Suslin, and the author, we further investigate refined invariants for finite dimensional representations $M$ of infinitesimal group schemes $\mathbb{G}$ over a field $k$ of characteristic $p>0$. In particular, we explore invariants of modules for Frobenius kernels $\mathbb{G}_{(r)}$ of linear algebraic groups $\mathbb{G}$ of exponential type. Such finite group schemes admit a refinement of support theory, including Jordan type functions and constructions of vector bundles for special classes of $\mathbb{G}_{(r)}$-modules. We reformulate earlier results for $\mathbb{G}_{(r)}$-modules in terms which are both more explicit and more accessible to computation, yet preserve the usual support theory for $\mathbb{G}_{(r)}$-modules.
@@
Jordan type, infinitesimal group scheme, vector bundles
@@ 0. Introduction
In the present work, we utilize methods and results of support theory to add to our understanding of invariants finer than support theory for certain classes of infinitesimal group schemes over an arbitrary base field $k$ of characteristic $p$. This is in contrast to efforts devoted to establish properties of one or more categories of representations of such a finite group scheme. Our techniques do not apply to finite groups, for we utilize the support theory developed in collaboration with A. Suslin and C. Bendel in [22], [23] which replaces cohomological varieties by varieties of 1-parameter subgroups.
@@
In particular, we investigate the “local Jordan type function” as briefly considered by J. Pevtsova and the author in [13, Defn 2.3]. For a given affine group scheme $\mathbb{G}$ of finite type over $k$ and a finite dimensional $\mathbb{G}$-module $M$ for the $r$-th Frobenius kernel $\mathbb{G}_{(r)}$ of $\mathbb{G}$, we formulate the continuous function  $JT_{\mathbb{G},r,M}(-):\mathbb{P}V_{r}(\mathbb{G})\to\mathcal{Y}$ from the scheme-theoretic points of $\mathbb{P}V_{r}(\mathbb{G})$ representing height $r$ 1-parameter subgroups to the partially ordered set $\mathcal{Y}$ of Young diagrams with $p$ columns. This “theory” $M\mapsto JT_{\mathbb{G},r,M}(-)$ captures much more information about a $\mathbb{G}_{(r)}$-module $M$ than does the support theory $M\mapsto\mathbb{P}V_{r}(\mathbb{G})_{M}$, yet possesses many good properties satisified by the more familiar cohomological support theory $M\mapsto\mathbb{P}H^{*}(\mathbb{G}_{(r)},k)_{M}$. For example, if $p$ does not divide the dimension of $M$, then $\mathbb{P}H^{*}(\mathbb{G}_{(r)},k)_{M}$ equals $\mathbb{P}H^{*}(\mathbb{G}_{(r)},k)$ and therefore is independent of $M$, whereas $M\ \mapsto\ JT_{\mathbb{G},r,M}(-)$ distinguishes many such $\mathbb{G}_{(r)}$-modules $M$ of a given dimension.
As a prelude to establishing the continuity of $JT_{\mathbb{G},r,M}(-)$, we reprove from first principles the compatibility of specialization of $M$ at a scheme-theoretic point of $\mathbb{P}V_{r}(\mathbb{G})$ with specialization of the Jordan type of $M$ at that scheme-theoretic point. A central role in this present work is played by the “universal $p$-nilpotent operator” $\Theta_{\mathbb{G},r}\in k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{(r)}$ constructed by Pevtsova and the author in [14]. Our first use of $\Theta_{\mathbb{G},r}$ is to complete the proof of the continuity of the function $JT_{\mathbb{G},r,M}(-)$ using the “projective extension” $\tilde{\Theta}_{\mathbb{G},r}$ of $\Theta_{\mathbb{G},r}$.
@@
In order to make the Jordan type function more explicit and computationally accessible, we restrict our attention to affine group schemes $\mathbb{G}$ which are of exponential type of height $r$; for such $\mathbb{G}$, height $r$ 1-parameter subgroups of $\mathbb{G}$ are parametrized by the variety $\mathcal{C}_{r}(\mathfrak{g})$ of $r$-tuples of $p$-nilpotent, pair-wise commuting elements of the Lie algebra $\mathfrak{g}$ of $\mathbb{G}$. This enables us to work with the explicitly defined variety $\mathcal{C}_{r}(\mathfrak{g})$ rather than the scheme $V_{r}(\mathbb{G})$ which is defined by the functor it represents.
For a given finite dimensional $\mathbb{G}_{(r)}$-module $M$, the Jordan type function becomes a continuous function $JT_{\mathfrak{g},r,M}(-):\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})\to\mathcal{Y}$ associated to the “projectivization” $\tilde{\Theta}_{\mathfrak{g},r}$ of the $p$-nilpotent element in $\Theta_{\mathfrak{g},r}\in k[\mathcal{C}_{r}(\mathfrak{g}]\otimes k\mathbb{G}_{(r)}$ corresponding to $\Theta_{\mathbb{G},r}$.
@@
We proceed to simplify this construction, replacing the function $JT_{\mathfrak{g},r,M}(-)$ by its linearization $JT_{\mathfrak{g},r,M}^{exp}(-):\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})\to\mathcal{Y}$ defined by replacing $\Theta_{\mathfrak{g},r}$ by its linear part (denoted $\Theta_{\mathfrak{g},r}^{exp}$) with respect to the grading on $\mathcal{C}_{r}(\mathfrak{g})$ transferred from the natural grading on $V_{r}(\mathbb{G})$. In the very special case that $\mathbb{G}$ equals $\mathbb{G}_{a(r)}$, this simplification is closely related to J. Carlson’s use of cyclic shifted subgroups to identify support varieties for $(\mathbb{Z}/p)^{\times r}$-modules (see [3]). Unlike Carlson’s cyclic shifted subgroups for $(\mathbb{Z}/p)^{\times r}$, $\Theta_{\mathfrak{g}_{a},r}^{exp}$ does not require a choice of generators for the augmentation ideal of $k\mathbb{G}_{a(r)}$.
@@
The suitability of $\Theta_{\mathfrak{g},r}^{exp}$ follows from work of Pevtsova, Suslin, and the author [15] and subsequent work of P. Sobaje [20]. The replacement of $JT_{\mathbb{G},r,M}(-)$ by $JT_{\mathfrak{g},r,M}^{exp}(-):\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})\to\mathcal{Y}$ does not affect certain refined support varieties considered in previous work of Pevtsova and the author in [13] and [14]. Moreover, we verify that vector bundles associated to $\mathbb{G}_{(r)}$-modules $M$ of constant Jordan type using $\tilde{\Theta}_{\mathfrak{g},r}^{exp}$ (the projectivization of $\Theta_{\mathfrak{g},r}^{exp}$) are $\mathbb{A}^{1}$-equivalent to vector bundles previously constructed in [14] using $\tilde{\Theta}_{\mathbb{G},r}$.
We provide sample computations which suggest the accessibility of $JT_{\mathfrak{g},r,M}^{exp}(-)$ for certain $\mathbb{G}_{(r)}$-modules. We also include some relevant discussion of weighted projective varieties and of the connection between 1-parameter subgroups for $\mathbb{G}$ and operators on $\mathbb{G}$-modules.
@@ 1. Jordan types and specialization
Let $M$ be a $k[t]/t^{p}$-module of dimension $m$ given by the map $\rho_{M}:k[t]/t^{p}\to M_{m,m}(k)$ of $k$-algebras. Let $\underline{a}=(a_{1},\ldots,a_{p})\in\mathbb{N}^{\times p}$ be defined by setting $a_{i}$ to be the number of blocks of size $i$ of the Jordan canonical form of the matrix $\rho_{M}(t)$. We say that the Jordan type of $M$ is $\underline{a}$, expressing this as $JT(M)=\underline{a}$.
@@
Let $M$ be a $k[t]/t^{p}$-module of dimension $m$ such that $JT(M)=\underline{a}$ (often written $JT(M)=a_{p}[p]+a_{p-1}[p-1]+\cdots+a_{1}[1]$).
@@
    The usual partial ordering of partitions determines the partial ordering on ×$\mathbb{N}^{\times p}$ given by
@@
    The Jordan type of $t^{j}$ acting on $M\simeq\oplus_{i=1}^{p}(k[t]/t^{i})^{\oplus a_{i}}$ (denoted by $\underline{a}^{j}\in\mathbb{N}^{\times p}$) has associated Young diagram obtained from the Young diagram for $\underline{a}$ by removing the first (i.e., left-most) $j-1$ columns of $\underline{a}$ and adding rows with each having only 1 box so that the number of boxes of $\underline{a}^{j}$ equals $m$.
@@
[…]
@@
Let ×$\mathbb{N}^{\times p}$ denote the monoid of $p$-tuples of non-negative integers $\underline{a}=(a_{1},\ldots,a_{p})$. Denote by $\mathcal{Y}_{\leq}$ the monoid ×$\mathbb{N}^{\times p}$ provided with the partial ordering of Proposition 1.2(3), and denote by $\mathcal{Y}_{\leq,n}$ the partially ordered subset of $\mathcal{Y}_{\leq}$ whose objects are $p$-tuples $\underline{a}$ with $\sum_{i=1}^{p}a_{i}\cdot i$ equal to $n$.
@@
of $k[t]t^{p}$-modules (see [5, Cor10.3]).
We recall the lower semi-continuity of Jordan type, an easy consequence of Nakayama’s Lemma (see [15, Prop 3.2]).
@@
Consider finitely generated field extensions $k\subset K\subset L$ and a noetherian, local domain $R\subset L$ with field of fractions $L$ and residue field $K$. Let $M_{R}$ be a free $R$-module of finite rank equipped with a $p$-nilpotent endomorphism $\theta_{R}\in End_{R}(M_{R},M_{R})$. Denote $K\otimes_{R}M$ by $M_{K}$ and $L\otimes_{R}M$ by $M_{L}$, and set
@@
By applying Nakayma’s Lemma [19, Lem4.3] to the cokernels of the maps $(\theta_{R})^{i}:M_{R}\to M_{R}$, we conclude that $dim(coker(\theta_{L}^{i}))\leq dim(coker(\theta_{K}^{i}))$ so that $rk(\theta_{K}^{i})\leq rk(\theta_{L}^{i})$ for each $i,1\leq i<p$. Thus, the proposition follows from Proposition 1.3. ∎
@@
Let $X$ be a scheme of finite type over $k$. For two scheme-theoretic points $x,y$ of $X$, we write $x\leq y$ if $x$ is in the (Zariski) closure of $y$. If $\operatorname{Spec}\nolimits R\subset X$ is an affine open subset of $X$ containing $x$, then the scheme-theoretic point $x\in\operatorname{Spec}\nolimits R$ corresponds to the prime ideal $\mathfrak{p}_{x}\subset R$. The structure sheaf $\mathcal{O}_{X}$ of $X$ has stalk the local ring $R_{(\mathfrak{p}_{x})}$ with residue field $k(x)=R_{(\mathfrak{p}_{x})}/\mathfrak{p}_{x}R_{(\mathfrak{p}_{x})}$. We identify the scheme-theoretic point $x\in X$ with the associated map of schemes $\chi_{x}:\operatorname{Spec}\nolimits k(x)\to X$.
We define $X_{\leq}$ to be the set of scheme-theoretic points of $X$ with this partial ordering.
@@
The following proposition will be employed with $A=k[V_{r}(\mathbb{G})]$ as in Definition 2.1 and with the action of $t$ given by various powers of the universal $p$-nilpotent operator of Theorem 2.11. We point out that in this proposition we consider the tensor product over the commutative algebra $A$ of $A$-modules; in discussion below, we shall frequently consider tensor product over $k$ of $\mathbb{G}$-modules using the Hopf algebra structure of $[\mathbb{G}]$.
@@
Let $A$ be a Noetherian commutative $k$-algebra and consider a finitely generated, projective $A$-module $M_{A}$ equipped with a $p$-nilpotent endomorphism $\theta_{A}\in Hom_{A}(M_{A},M_{A})$. Then sending a scheme-theoretic point $x\in\operatorname{Spec}\nolimits A$ (i.e., a prime ideal $\mathfrak{p}_{x}\subset A$) to the Jordan type of $k(x)\otimes_{A}M_{A}$ as a $k(x)[t]/t^{p}$-module determines a map of posets
(1.9.1) $JT_{A,M_{A}}(-):(\operatorname{Spec}\nolimits A)_{\leq}\quad\to\quad\mathcal{Y}_{\leq},\quad x\mapsto JT(k(x)\otimes_{A}M_{A}).$
@@
$JT_{A,M_{A}}(x)\oplus JT_{A,N_{A}}(x)\ =\ JT_{A,M_{A}\oplus N_{A}}(x),\quad JT_{A,M_{A}}(x)\otimes JT_{N_{A}}(x)\ =\ JT_{A,M_{A}\otimes_{A}N_{S}}(x).$
@@
The assertion that $JT_{A,M_{A}}(-)$ is a map of posets is the assertion that $x\leq y\in\operatorname{Spec}\nolimits A$ implies that $JT_{A,M_{A}}(x)\leq JT_{N,N_{A}}(y)$. As in Lemma 1.8, we may restrict $M_{A}$ along a map from $A$ to $R$ with a local domain with field of fractions $k(y)$ and residue field $k(x)$. Thus, we reduce to the case that $A=R$ is a local domain with field of fractions $k(y)$ and residue field $k(x)$. In this case, Proposition 1.6 applies to prove $JT_{A,M_{A}}(x)\leq JT_{A,M_{A}}(y)$.
The fact that the function $JT_{A,-}(x)$ (sending a finitely generated $A[t]/t^{p}$-module $M_{A}$ to $JT_{A,M_{A}}(x)\in\mathcal{Y}_{\leq}$) respects both addition and tensor product is immediate from the definitions of addition and tensor product for $\mathcal{Y}_{\leq}$. ∎
Proposition 1.9 does not immediately imply that $JT_{A,M_{A}}(-):\operatorname{Spec}\nolimits A\ \to\ \mathcal{Y}_{\leq,m}$ is continuous when $\operatorname{Spec}\nolimits A$ is given the Zariski topology. This is because arbitrary unions of closed subsets (i.e., subsets closed under specialization) are often not closed in the Zariski topology. We supplement Proposition 1.9 with another consequence of Nakayama’s Lemma in commutative algebra. (It seems more convenient for this lemma to refer to the scheme-theoretic points of $\operatorname{Spec}\nolimits A$ as prime ideals $\mathfrak{p}\subset A$.)
@@
We remind the reader that an Alexandrov discrete topology on a $T_{0}$ space is a topology in which the arbitrary union of closed subsets is closed. (See [1] and [18].) The topology on $\mathcal{Y}$ which we utilize in the following theorem is such a topology.

### v2

@@ Abstract.
We revisit the constructions given by J. Pevtsova and the author of refined invariants for finite dimensional representations of infinitesimal group schemes $\mathbb{G}_{(r)}$ over a field $k$ of characteristic $p>0$. Our focus is on the universal $p$-nilpotent operator seen as an element in the group algebra of the group scheme $\mathbb{G}_{(r),X}$ over $X$, where $X$ is either the moduli space $V_{r}(\mathbb{G})$ of height $r$ 11-parameter subgroups of $\mathbb{G}$ or the moduli space $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ of $r$-tuples of $p$-nilpotent, pair-wise commuting elements of the Lie algebra of $\mathbb{G}$.
We formalize Jordan type function using several variants of the continuous function $JT_{\mathbb{G},r,M}(-):\mathbb{P}V_{r}(\mathbb{G})\to\mathcal{Y}$ where $\mathcal{Y}$ is the poset of Young diagrams with $p$-columns. One of these variants is designed to be more conducive to computation. The vector bundle construction given by J. Pevtsova and the author is extended to all finite dimensional $\mathbb{G}_{(r)}$-modules, producing coherent sheaves on $X$ which are locally free on the strata of $X$ associated to $JT_{\mathbb{G},r,M}(-)$.
@@
Jordan type, infinitesimal group scheme, universal operators
@@ 0. Introduction
In the present work, we utilize methods and results of support theory to investigate explicit invariants for specific finite dimensional modules for certain classes of infinitesimal group schemes over an arbitrary base field $k$ of characteristic $p$. This is in contrast to the role of support theory in establishing properties for various categories of representations of finite group schemes as in [3]. Our techniques do not apply to finite groups, for we utilize the support theory developed in collaboration with A. Suslin and C. Bendel in [26], [27] which replaces cohomological varieties by varieties of 1-parameter subgroups.
@@
We further investigate the “Jordan type function” introduced by J. Pevtsova and the author in [15] and [16]. For a given affine group scheme $\mathbb{G}$ of finite type over $k$ and a finite dimensional $\mathbb{G}$-module $M$ for the $r$-th Frobenius kernel $\mathbb{G}_{(r)}$ of $\mathbb{G}$, we formulate the continuous function $JT_{\mathbb{G},r,M}(-):\mathbb{P}V_{r}(\mathbb{G})\to\mathcal{Y}$ from the scheme-theoretic points of $\mathbb{P}V_{r}(\mathbb{G})$ representing height $r$ 1-parameter subgroups to the partially ordered set $\mathcal{Y}$ of Young diagrams with $p$ columns. By definition, $JT_{\mathbb{G},r,M}(-)=JT_{\mathbb{G}_{(r)},r,M}(-)$. The continuity of $JT_{\mathbb{G},r,M}(-)$ implies that if a point $x\in\mathbb{P}V_{r}(\mathbb{G})$ specializes to $y\in\mathbb{P}V_{r}(\mathbb{G})$ then the Young diagram $JT_{\mathbb{G},r,M}(x)$ is greater or equal to the Young diagram $JT_{\mathbb{G},r,M}(y)$.
This “theory” $M\mapsto JT_{\mathbb{G},r,M}(-)$ constitutes a functor $JT_{\mathbb{G},r}:mod(\mathbb{G}_{(r)})\to Hom_{cont}(\mathbb{P}V_{r}(\mathbb{G}),\mathcal{Y})$ from finite dimensional $\mathbb{G}_{(r)}$-modules to continuous functions $\mathbb{P}V_{r}(\mathbb{G})\to\mathcal{Y}$ which captures much more information about a $\mathbb{G}_{(r)}$-module $M$ than does the cohomological support theory $M\mapsto\mathbb{P}H^{*}(\mathbb{G}_{(r)},k)_{M}$. For example, if $p$ does not divide the dimension of $M$, then $\mathbb{P}H^{*}(\mathbb{G}_{(r)},k)_{M}$ equals $\mathbb{P}H^{*}(\mathbb{G}_{(r)},k)$ and therefore is independent of $M$, whereas $M\ \mapsto\ JT_{\mathbb{G},r,M}(-)$ distinguishes many such $\mathbb{G}_{(r)}$-modules $M$ of a given dimension. As might be expected, determination of $JT_{\mathbb{G},r,M}(-)$ is a challenge. With this in mind, we introduce a simplified Jordan type function $M\ \mapsto\ JT^{exp}_{\mathfrak{g},r,M}(-)$ for affine group schemes $\mathbb{G}$ which are of exponential type of height $r$.
After an elementary discussion of Jordan types, we provide a detailed formalism for the construction of the “universal $p$-nilpotent operator” $\Theta_{\mathbb{G},r}\in k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{(r)}\equiv k\mathbb{G}_{(r),V_{r}(\mathbb{G})}$ of height $r$ constructed by J. Pevtsova and the author in [16]. In doing so, we expand upon the connection between 1-parameter subgroups for $\mathbb{G}$ and operators on $\mathbb{G}_{(r)}$-modules. The operator $\Theta_{\mathbb{G},r}$ acting on $k[V_{r}(\mathbb{G})]\otimes M$ determines the Jordan type function $JT^{exp}_{\mathfrak{g},r,M}(-)$. The categorically inclined reader might want to categorify the functor $JT_{\mathbb{G},r}$ by replacing $Hom_{cont}(\mathbb{P}V_{r}(\mathbb{G}),\mathcal{Y})$ by the category of graded $k[V_{r}(\mathbb{G})]$-modules equipped with a $p$-nilpotent homogeneous operator.
@@
In order to make the Jordan type function more explicit and computationally accessible, we restrict our attention to affine group schemes $\mathbb{G}$ which are of exponential type of height $r$. By definition, such groups $\mathbb{G}$ are equipped with an exponential map with the property that height $r$ 1-parameter subgroups of $\mathbb{G}$ are given by applying the exponential map to $r$-tuples of $p$-nilpotent, pair-wise commuting elements of the Lie algebra $\mathfrak{g}$ of $\mathbb{G}$. Thus, the “moduli space” for height $r$ 1-parameter subgroups of $\mathbb{G}$ (of exponential type $r$) is the variety $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ of such $r$-tuples of elements in $\mathfrak{g}$. For $\mathbb{G}=GL_{N}$, the associated exponential map is the usual (truncated) exponential sending a $p$-nilpotent element of $A\otimes\mathfrak{g}l_{N}$ to its exponential in $GL_{N}(A)$ for any commutative $k$-algebra AA.
@@
For a given finite dimensional $\mathbb{G}_{(r)}$-module $M$ with $\mathbb{G}$ of exponential type of height $r$, the Jordan type function becomes a continuous function $JT_{\mathfrak{g},r,M}(-):\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\to\mathcal{Y}$ associated to the “projectivization” of the universal $p$-nilpotent operator $\Theta_{\mathfrak{g},r}\in k[\mathcal{C}_{r}(\mathcal{N}_{p}\mathfrak{g}))]\otimes k\mathbb{G}_{(r)}$ corresponding to $\Theta_{\mathbb{G},r}$. Our exponential Jordan type function $JT_{\mathfrak{g},r,M}^{exp}(-):\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\to\mathcal{Y}$ is defined by replacing $\Theta_{\mathfrak{g},r}$ by its linear part (denoted $\Theta_{\mathfrak{g},r}^{exp}$) with respect to a natural grading on the group algebra $k\mathbb{G}_{a(r)}$ of $\mathbb{G}_{a(r)}$. In the very special case that $\mathbb{G}$ equals $\mathbb{G}_{a(r)}$, this simplification is related to J. Carlson’s use of cyclic shifted subgroups to identify support varieties for $(\mathbb{Z}/p)^{\times r}$-modules (see [5]).
@@
The suitability of $\Theta_{\mathfrak{g},r}^{exp}$ follows from work of Pevtsova, Suslin, and the author [17] and subsequent work of P. Sobaje [24]. Namely, for a $\mathbb{G}_{(r)}$-module $M$ of dimension $m$ and a point of $x\in\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g})$), $JT_{\mathfrak{g},r,M}(x)\ =\ m/p\cdot[p]$ if and only if $JT^{exp}_{\mathfrak{g},r,M}(x)\ =\ m/p\cdot[p]$. Thus, the replacement of $JT_{\mathbb{G},r,M}(-)$ by $JT_{\mathfrak{g},r,M}^{exp}(-):\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\to\mathcal{Y}$ does not affect (refined) support varieties considered in previous work of Pevtsova and the author in [15] and [16]. We provide sample computations which demonstrate the accessibility of $JT_{\mathfrak{g},r,M}^{exp}(-)$ for certain $\mathbb{G}_{(r)}$-modules $M$.
For $\mathbb{G}$ of exponential type of height $r$, we verify that vector bundles on $\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ associated to $\mathbb{G}_{(r)}$-modules $M$ of constant Jordan type using $\Theta_{\mathfrak{g},r}$ (naturally corresponding to vector bundles previously constructed in [16]) are $\mathbb{A}^{1}$-homotopy equivalent to vector bundles using $\Theta_{\mathfrak{g},r}^{exp}$. We extend these constructions of vector bundles to apply to arbitrary finite dimensional $\mathbb{G}_{(r)}$-modules $M$, yielding coherent sheaves on $\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ whose restrictions on strata associated to the Jordan type function for $M$ are locally free. Unlike the situation for modules of constant Jordan type, these stratified coherent sheaves associated to a given finite dimensional $\mathbb{G}_{(r)}$-module will typically depend upon whether we use the Jordan type function associated to $\Theta_{\mathbb{G},r}$ or associated to $\Theta_{\mathfrak{g},r}^{exp}$.
As we observe in the last section, the issue of “stabilization with respect to rr” of Jordan type functions of height $\leq r$ is closely related to the Jordan type function for finite dimensional rational $\mathbb{G}$-modules introduced in [12] if $\mathbb{G}$ is of exponential type of infinite height. This corrects a minor error of [12, Prop 4.8] which incorrectly formulated such a stabilization result.
@@ 1. Jordan types and specialization
Let $M$ be a $k[t]/t^{p}$-module of dimension $m$ given by the map $\rho_{M}:k[t]/t^{p}\to M_{m,m}(k)$ of $k$-algebras. Let $\underline{a}=(a_{1},\ldots,a_{p})\in\mathbb{N}^{\times p}$ (i.e., a $p$-tuple of non-negative integers) be defined by setting $a_{i}$ to be the number of blocks of size $i$ of the Jordan canonical form of the matrix $\rho_{M}(t)$. We say that the Jordan type of $M$ is $\underline{a}$, expressing this as $JT(M)=\underline{a}$ or as $JT(M)=a_{p}[p]+a_{p-1}[p-1]+\cdots+a_{1}[1]$.
@@
Let $M$ be a $k[t]/t^{p}$-module of dimension $m$ such that $JT(M)=\underline{a}$.
@@
    The usual partial ordering of partitions determines the partial ordering on $\mathbb{N}^{\times p}$ given by
@@
    The Jordan type of $t^{j}$ acting on $M\simeq\oplus_{i=1}^{p}(k[t]/t^{i})^{\oplus a_{i}}$ (denoted by $\underline{a}^{j}\in\mathbb{N}^{\times p}$) has associated Young diagram obtained from the Young diagram for $\underline{a}$ by removing the first (i.e., left-most) j1j-1 columns of $\underline{a}$ and adding rows with each having only 1 box so that the number of boxes of $\underline{a}^{j}$ equals $m$.
@@
Consider $m$-dimensional $k[t]/t^{p}$-modules M,NM,\ N isomorphic to $\bigoplus_{i=1}^{p}(k[t]/t^{i})^{\oplus a_{i}},\ \bigoplus_{i=1}^{p}(k[t]/t^{i})^{\oplus b_{i}}$. As above, let $\underline{a}^{s}$ (respectively, $\underline{b}^{s}$) denotes the Jordan type of $t^{s}$ acting on $M$ (resp, $N$) for any $s,\ 1\leq s<p$.
@@
Denote by $\mathcal{Y}_{\leq}$ the monoid $\mathbb{N}^{\times p}$ provided with the partial ordering of Proposition 1.2(3), and denote by $\mathcal{Y}_{\leq,n}$ the partially ordered subset of $\mathcal{Y}_{\leq}$ whose objects are $p$-tuples $\underline{a}$ with $\sum_{i=1}^{p}a_{i}\cdot\ \leq\ n$.
@@
of $k[t]t^{p}$-modules (see [7, Cor10.3]).
We recall the lower semi-continuity of Jordan type, an easy consequence of Nakayama’s Lemma (see [17, Prop 3.2]).
@@
Consider a noetherian, local domain $R$ with field of fractions $L$ and residue field $K$. Let $M_{R}$ be a free $R$-module of finite rank equipped with a $p$-nilpotent endomorphism $\theta_{R}\in End_{R}(M_{R},M_{R})$. Denote $K\otimes_{R}M$ by $M_{K}$ and $L\otimes_{R}M$ by $M_{L}$, and set
@@
By applying Nakayma’s Lemma [22, Lem 4.3] to the cokernels of the maps $(\theta_{R})^{i}:M_{R}\to M_{R}$, we conclude that $dim(coker(\theta_{L}^{i}))\leq dim(coker(\theta_{K}^{i}))$ so that $rk(\theta_{K}^{i})\leq rk(\theta_{L}^{i})$ for each $i,\ 1\leq i<p$. Thus, the proposition follows from Proposition 1.3. ∎
@@
Let $X$ be a scheme of finite type over $k$. For two scheme-theoretic points $x,y$ of $X$, we write $x\leq y$ if $x$ is in the (Zariski) closure of $y$. This endows the set of scheme-theoretic points of $X$ with a natural partial ordering.
If $\operatorname{Spec}\nolimits R\subset X$ is an affine open subset of $X$ containing $x$, then the scheme-theoretic point $x\in\operatorname{Spec}\nolimits R$ corresponds to the prime ideal $\mathfrak{p}_{x}\subset R$. The structure sheaf $\mathcal{O}_{X}$ of $X$ has stalk the local ring $R_{(\mathfrak{p}_{x})}$ with residue field $k(x)=R_{(\mathfrak{p}_{x})}/\mathfrak{p}_{x}R_{(\mathfrak{p}_{x})}$. We identify the scheme-theoretic point $x\in X$ with the associated map of schemes $\chi_{x}:\operatorname{Spec}\nolimits k(x)\to X$.
@@
The following proposition will be employed with $A=k[V_{r}(\mathbb{G})]$ as in Definition 2.1 and with the action of $t$ given by various powers of the universal $p$-nilpotent operator of Theorem 2.15. We point out that in this proposition we consider the tensor product over the commutative algebra $A$ of $A$-modules; in discussion below, we shall frequently consider tensor product over $k$ of $\mathbb{G}$-modules using the Hopf algebra structure of $k[\mathbb{G}]$.
@@
Let $A$ be a Noetherian commutative $k$-algebra and consider a finitely generated, projective $A$-module $M_{A}$ equipped with a $p$-nilpotent endomorphism $\theta_{A}\in Hom_{A}(M_{A},M_{A})$. Then sending a scheme-theoretic point $x\in\operatorname{Spec}\nolimits A$ (i.e., a prime ideal $\mathfrak{p}_{x}\subset A$) to the Jordan type of $k(x)\otimes_{A}M_{A}$ as a $k(x)[t]/t^{p}$-module with $t$ acting as $k(x)\otimes\Theta_{A}\in Hom_{k(x)}(M\otimes k(x),M\otimes k(x))$ determines a map of partially ordered sets
(1.9.1) $JT_{A,M_{A}}(-):(\operatorname{Spec}\nolimits A)\quad\to\quad\mathcal{Y}_{\leq}\ ,\quad x\mapsto JT(k(x)\otimes_{A}M_{A}).$
@@
$JT_{A,M_{A}}(x)\oplus JT_{A,N_{A}}(x)\quad=\quad JT_{A,M_{A}\oplus N_{A}}(x),$
$JT_{A,M_{A}}(x)\otimes JT_{A,N_{A}}(x)\quad=\quad JT_{A,M_{A}\otimes_{A}N_{S}}(x).$
@@
The assertion that $JT_{A,M_{A}}(-)$ is a map of posets is the assertion that $x\leq y\in\operatorname{Spec}\nolimits A$ implies that $JT_{A,M_{A}}(x)\leq JT_{N,N_{A}}(y)$. As in Lemma 1.8, we may restrict $M_{A}$ along a map from $A$ to a local domain $R$ with field of fractions $k(y)$ and residue field $k(x)$. Thus, we reduce to the case that $A=R$ is a local domain with field of fractions $k(y)$ and residue field $k(x)$; in this case, $M_{A}$ is a free $A$-module of finite rank. Thus, Proposition 1.6 applies to prove $JT_{A,M_{A}}(x)\leq JT_{A,M_{A}}(y)$.
Observe that specialization along a scheme-theoretic point of $\operatorname{Spec}\nolimits A$ commutes with direct sums and tensor products. Thus, the fact that the function $JT_{A,-}(x)$ (sending a finitely generated $A[t]/t^{p}$-module $M_{A}$ to $JT_{A,M_{A}}(x)\in\mathcal{Y}_{\leq}$) respects both addition and tensor product is immediate from the definitions of addition and tensor product for $\mathcal{Y}_{\leq}$. ∎
Proposition 1.9 does not immediately imply that $JT_{A,M_{A}}(-):\operatorname{Spec}\nolimits A\ \to\ \mathcal{Y}_{\leq}$ is continuous when $\operatorname{Spec}\nolimits A$ is given the Zariski topology. This is because arbitrary unions of closed subsets (i.e., subsets closed under specialization) are often not closed in the Zariski topology. We supplement Proposition 1.9 with another consequence of Nakayama’s Lemma in commutative algebra. (It seems more convenient for this lemma to refer to the scheme-theoretic points of $\operatorname{Spec}\nolimits A$ as prime ideals $\mathfrak{p}\subset A$.)
@@
We remind the reader that an Alexandrov discrete topology on a $T_{0}$ space is a topology in which the arbitrary union of closed subsets is closed. (See [1] and [21].) The topology on $\mathcal{Y}_{\leq}$ which we utilize is such a topology.
Definition 1.11.
We give $\mathcal{Y}_{\leq}$ the topology whose closed subsets are those subsets $S\subset\mathcal{Y}_{\leq}$ satisfying the condition that if $s\in S$ and $s^{\prime}\leq s$ then $s^{\prime}\in S$.
In particular, $S\subset\mathcal{Y}_{\leq}$ is closed if and only if $S\cap\mathcal{Y}_{\leq,m}$ is closed in $\mathcal{Y}_{\leq,m}$ for all $m>0$.